At what velocity shall a dropped object fall a distance of 1.75 m? Neglect air resistance.


1
Expert's answer
2021-11-15T10:33:12-0500

"H=\large\frac{v-v_0}{2g}\to" "v_o=0 \to v = \sqrt{2gH}=\sqrt{2*10*1.75}=\sqrt{35}=5.916 \frac{m}{s} \thickapprox 6 \frac{m}{s}"